OpenStudy (anonymous):

A railroad track and a road cross at right angles. An observer stands on the road 70 meters south of the crossing and watches an eastbound train traveling at 60 meters per second. At how many meters per second is the train moving away from the observer 4 seconds after it passes through the intersection?

OpenStudy (rulnick):

The component of the train's velocity in the direction away from the observer is 60cos(atan(70/(60*4))) = 60cos(atan(7/24)) = 60(0.96) = 57.6 meters per second.